Circular No. 8255

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 104P/KOWAL On 1973 Jan. 14, the Central Bureau received a cablegram with the

following two-line message from the late L. Boethin, Abra, Philippines (cf. IAUC 2745): “new comet january 11 20:00 ut in 11h 38m” and “20s 9m 5 moving sse 8 minutes a day”. Although there was uncertainty as to whether the “20s” referred to the seconds of time of α or to the fact that δ was “20^{o} south” (for otherwise the δ information was absent), visual searches were conducted at the request of the Bureau to mag 13 over the general range 11^{h} ≤ α ≤ 12^{h}, +3^{o} ≥ δ ≥ −30^{o}. The searches were unsuccessful, and Boethin did not respond to a cabled request that he verify the Bureau’s interpretation of his cablegram. A letter received by the Bureau on Jan. 29 showed that “20s” did refer to the seconds of α but that, by chance, δ really was −20^{o}; the magnitude was given as 9.5, so that, if the comet existed, the confirmation attempts ought to have revealed it. The letter also mentioned that the coma was large and diffuse, 8′ in diameter, with a small distinct nucleus. It included the following three positions (made visually with a 0.20-m reflector and converted here to equinox 2000.0) and was dated and timed coincidentally with the third observation:

1973 UT  α_{2000}  δ_{2000}
Jan. 11.833  11^{h}40^{m}52^{s}  −20^{o}29′
12.833  11 41 01  −20 37
13.833  11 41 10  −20 45

In a special-delivery letter dated Feb. 11, Boethin added that he had observed the comet again “on Monday” (presumably on Jan. 14.8 UT) but that it had by then “darkened considerably”, to “about mag 12–13!”. “On Tuesday” the bright moon (it was full on Jan. 18.9 UT) prevented further observations. G. Kronk, Troy, IL, who was in communication with Boethin in the 1980s concerning the latter’s several unconfirmed comets (Marsden and Roemer 1978, Quart. J. Roy. Astron. Soc. 19, 64), has recently suggested that the 1973 object might have been comet 104P. At that time, some months after the perihelion passage (T = 1972 Aug. 4) preceding that of its 1979 discovery, comet 104P would not have been expected to be brighter than mag 17. However, the nongravitational forces affecting this comet are unusually large and variable (indeed, it was necessary to incorporate them in a linkage of the first two apparitions, as well as for the 1997–2003 linkage), a circumstance well known to be associated with outbursts (and splitting). Nevertheless, a nongravitational solution reasonably fitting the 1979–1998 observations does represent Boethin’s three rough positions within 2′, so the identification is secure.

2003 December 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Brian G. Marsden